A high school graduating class is made up of 514 students. There are 108 more girls than boys. How many boys are in the class?
Answer by algebrahouse.com(1659)

"A high school graduating class is made up of 514 students. There are 108 more girls than boys. How many boys are in the class?"

x = number of boys
x + 108 = number of girls {there are 108 more girls than boys}

x + x + 108 = 514 {there are 514 total students}
2x + 108 = 514 {combined like terms}
2x = 406 {subtracted 108 from both sides}
x = 203 {divided both sides by 2}

203 boys in the class

For more help from me, visit: www.algebrahouse.com